# Gradient-Free Adversarial Purification with Diffusion Models

### v1

@@ 4.2.1 Adversarial Anti-Aliasing
Anti-aliasing is a straightforward, zero-shot method for smoothing image details, including adversarial perturbations [21, 32]. Unlike previous works, we have found that anti-aliasing with non-square filters is particularly effective against adversarial attacks while preserving clean accuracy. Additionally, using the average value from neighboring pixels, excluding the original pixel, has also proven effective. This is because adversarial perturbations are calculated on a pixel-wise basis and are sensitive to pixel changes. These two approaches greatly enhance the effectiveness of anti-aliasing. Even with simple anti-aliasing, we achieve moderate defense performance, underscoring the effectiveness of our approach. The detailed calculation method is discussed in the appendix. Although adversarial anti-aliasing can produce blurred images, the semantic features are preserved because the adversarial perturbation should remain imperceptible. Therefore, it effectively reduces the magnitude of adversarial perturbations while maintaining the semantic information necessary for classification. To maintain the resolution of the output image, we use padding, which is calculated as follows:
[…]
where $R$ is the shape of the data. We use stride $=1$.
@@
Table 1: The standard and robust accuracy against left: AutoAttack ($\ell_{\text{inf}}=8/255$), right: PGD-EOT ($\ell_{\text{inf}}=8/255$) on CIFAR10.
Method Target Model Standard Acc(%) AutoAttack Acc(%) PGD-EOT Acc(%)
@@
4.2.2 Adversarial Super-Resolution
During the adversarial anti-aliasing phase, we significantly reduce adversarial perturbations by directly decreasing pixel-wise modifications of adversarial examples. However, this approach may not be effective against unrestricted adversarial examples, as they are not generated by adding explicit perturbations. Additionally, blurring the images can negatively impact the clean accuracy of the target model. Super-resolution offers an effective way to recover high-quality images from our adversarial anti-aliased images. Previous super-resolution methods [19, 12] typically modify the original pixels of the low-resolution image and use the residual features of the original low-resolution image. These methods can inadvertently transfer negative effects from the adversarial examples to the final high-resolution images, making them ineffective for adversarial super-resolution. Diffusion-model-based super-resolution [38, 26] provides a more isolated approach for super-resolution. These models generate high-resolution images through a denoising-like process over randomly sampled noise, using the low-resolution image as conditions.
In this work, we adopt the ResShift method by Yue et al. [38] for our super-resolution process. This super-resolution model can also incorporate benign priors for defense, as it is trained with the clean dataset of the target model. Figure 4 demonstrates that the proposed super-resolution method achieves results comparable to diffusion-based adversarial purification [29], which do not require the calculation of gradient.
4.2.3 Adversarial Purification
$y=\{f(\text{SR}(\text{AA}(x_{\text{adv}}))))\}$ (9)
4.3 Discussions on Improved Time Efficiency
Second, we employ diffusion-based super-resolution instead of diffusion-based image generation. It is well-known that the reverse process of diffusion models is time-consuming, as illustrated in Figure 2, and the gradient calculation exacerbates this issue. However, we may not require the entire reverse process for purification, given that we already have a reference adversarial example, which is also discussed in Nie et al. [24]’s work. Since adversarial perturbations are pixel-wise, we opt for a relatively lightweight generation task, namely super-resolution, which also focuses on pixel modification. The diffusion-based super-resolution method used in this paper requires only tens of steps, compared to the hundreds or thousands of diffusion steps needed in previous works. With these two approaches, we significantly enhance the time efficiency of diffusion-based adversarial purification without compromising defense performance.
5 Experiments
5.1 Experimental Setup
Dataset and target models. We consider CIFAR-10 [18] and ImageNet [10] for major evaluation. For target models, we adopt WideResNet-28-10 and WideResNet-70-16 [39] for CIFAR-10 dataset and ResNet50 [14] for ImageNet dataset. These are commonly adopted backbones for adversarial robustness evaluation.
Attack settings. We evaluate our method with both perturbation-based attacks and diffusion-based unrestricted adversarial attacks. For perturbation-based attacks, we select AutoAttack [7], PGD [23]. For diffusion-based unrestricted adversarial attacks, we use DiffAttack [4] and AdvDiff [9] for comparisons. DiffAttack is only evaluated on the ImageNet dataset according to the original paper. To ensure a fair comparison with previous diffusion-based adversarial purification, we include the evaluation against the adaptive attack, i.e., reverse pass differentiable approximation (BPDA) [15]. We also evaluate the performance against PGD+EOT that is discussed in [20]. On CIFAR-10, the attack settings follow DiffPure [24]. On ImageNet, we randomly sample 5 images from each class and average over 10 runs. The PGD+EOT settings all follow Lee et al. [20].
Implementation details. We adopt the mean filter with $[[1,1],[1,1]]$ for adversarial anti-aliasing on CIFAR-10, and $[[1,1,1,1,1],[1,1,0,1,1],[1,1,1,1,1]]$ in ImageNet. ResShift [38] is utilized for adversarial super-resolution. We use the official Score SDE [31] checkpoint for CIFAR-10 and LDM [26] checkpoint for ImageNet to generate UAEs. More details and experiment results are given in the appendix.
Evaluation metrics. Following Nie et al. [24], we use standard accuracy and robust accuracy as the evaluation metrics. Both are calculated according to the top-1 classification accuracy.
5.2 Attack Performance
5.2.1 CIFAR10
Perturbation-based adversarial attack. Table 1 presents the defense performance against AutoAttack ($\ell_{\text{inf}}=8/255$) on the CIFAR10 dataset. The results demonstrate that our proposed method achieves better standard accuracy and robust accuracy than previous attack methods. Because images in the CIFAR10 dataset are only with $32\times 32$ resolution, we set our anti-aliasing filter to a relatively small size. Table 2 indicates that the robustness performance of the proposed method is on par with the state-of-the-art method [24]. This finding suggests that our method is more suitable for high-resolution images, as $32\times 32$ may not be large enough to effectively extract the semantic shape for our approach. However, we can further enhance our performance by incorporating adversarial purification techniques from previous work.
Our defense’s performance against PGD-EOT showcases its ability to defend against adaptive attacks. This is because our approach focuses on extracting and recovering the semantic features from adversarial images, rather than inferring and denoising the adversarial perturbations. As a result, our defense maintains similar effectiveness against both adaptive and standard attacks.
@@ 5.2.1 CIFAR10
Table 2: The standard and robust accuracy against BPDA ($\ell_{\text{inf}}=8/255$) on the CIFAR10 dataset with WideResNet-28-10 as the target model.
@@
Wang et al. [33] Diffusion 90.36 77.31
Song et al. [29] Diffusion 91.41 76.45
Ours Diffusion 91.52 $\pm$ 1.28 81.24 $\pm$ 2.51
Unrestricted adversarial attack. Unrestricted adversarial examples on the CIFAR10 dataset are challenging to defend against, as shown in Table 3. Our purification method outperforms the previous adversarial purification approach [29] by an average of 10%, validating the effectiveness of our proposed defense.
Table 3: The standard and robust accuracy against AdvDiff on the CIFAR10 dataset.
@@
Table 4: The standard and robust accuracy against AutoAttack ($\ell_{\text{inf}}=8/255$) on the ImageNet dataset.
Method Target Model Standard Acc(%) Robust Acc(%)
Engstrom et al. [8] ResNet50 62.56 31.06
Wong et al. [34] ResNet50 55.62 26.95
Salman et al. [27] ResNet50 64.02 37.89
Bai et al. [2] ResNet50 67.38 35.51
Nie et al. [24] ResNet50 68.22 43.89
Song et al. [29] ResNet50 66.92 61.53
Ours ResNet50 75.28 $\pm$ 1.06 67.61 $\pm$ 1.95
Table 5: The standard and robust accuracy against left: PGD ($\ell_{\text{inf}}=4/255$), right: PGD+EOT ($\ell_{\text{inf}}=4/255$) on ImageNet dataset.

### v2

@@ IV-B1 Adversarial Anti-Aliasing
Anti-aliasing is a straightforward, zero-shot method for smoothing image details, including adversarial perturbations [35, 36]. Unlike previous works, we have found that anti-aliasing with non-square filters is particularly effective against adversarial attacks while preserving clean accuracy. Additionally, using the average value from neighboring pixels, excluding the original pixel, has also proven effective. This is because adversarial perturbations are calculated on a pixel-wise basis and are sensitive to pixel changes. These two approaches greatly enhance the effectiveness of anti-aliasing. Even with simple anti-aliasing, we achieve moderate defense performance, underscoring the effectiveness of our approach. We give the details on the implementation of proposed adversarial anti-aliasing in Figure 4. We use 1D convolution on each RBG channel of the adversarial image and clip the additional dimensions according to the input resolution. After that, we convert the anti-aliased image tensor to the RGB image for adversarial super-resolution.
To maintain the resolution of the output image, we use padding, which is calculated as follows:
$R_{out}=\left\lfloor R_{in}+2\times\text{Padding}-\text{filter\_size}\right\rfloor$ (8)
where $R$ is the shape of the data. We use stride =1=1.
The proposed adversarial anti-aliasing utilizes the mean value from a pixel’s neighborhood and can be seen as a direct method for reducing adversarial perturbations. It is well-known in image classification that the areas contributing most to classification results are typically located at the center of the main object, where they are rich in textures. Consequently, the anti-aliasing filter can effectively neutralize adversarial perturbations in these regions, which often exhibit significant changes in RGB values and gradients of the classification loss function [37], especially in high-resolution images. By applying blurring, the proposed adversarial anti-aliasing method can significantly decrease the magnitude of adversarial perturbations. The worst case of the proposed method occurs in simple backgrounds, such as the sky. However, these areas generally have minimal impact on the final classification results.
IV-B2 Adversarial Super-Resolution
Figure 5: The example of proposed adversarial super-resolution. Our method achieves similar adversarial purification without any gradient calculation of diffusion models. It is worth noting that our adversarial super-resolution achieves superior restoration of background regions in the image.
During the adversarial anti-aliasing phase, we significantly reduce adversarial perturbations by directly decreasing pixel-wise modifications of adversarial examples. However, this approach may not be effective against unrestricted adversarial examples, as they are not generated by adding explicit perturbations. Additionally, blurring the images can negatively impact the clean accuracy of the target model. Super-resolution offers an effective way to recover high-quality images from our adversarial anti-aliased images. Previous super-resolution methods [38, 39] typically modify original pixels of the low-resolution image and use the residual features of the original low-resolution image. These methods can inadvertently transfer negative effects from the adversarial examples to the final high-resolution images, making them ineffective for adversarial super-resolution. Diffusion-based super-resolution [40, 10] provides a more isolated approach for super-resolution. These models generate high-resolution images through a denoising-like process over randomly sampled noise, using the low-resolution image as conditions.
In this work, we adopt the ResShift method by Yue et al. [40] for our super-resolution process. This super-resolution model can incorporate benign priors for defense, as it is trained with the clean dataset of the target model. ResShift is a diffusion-based model for image restoration, designed to learn the residual transformation $\bm{e}_{0}=\bm{x}_{0}-\bm{X}$ that maps a low-resolution (LR) image $\bm{x}_{0}$ to its corresponding high-resolution (HR) image $\bm{X}$. The residual $\bm{e}_{0}$ is scheduled over time steps $\eta_{t}$}, where $t=1,2,\dots,T$ denotes the diffusion timesteps. The forward process of ResShift is defined as:
$q(\bm{x}_{t}|\bm{x}_{t-1},\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\bm{x}_{t-1}+\alpha_{t}\bm{e}_{0},\kappa^{2}\alpha_{t}\textbf{I})$ (9)
where $\alpha_{t}=\eta_{t}-\eta_{t-1}$ for $t>1$ and $\alpha_{1}=\eta_{1}$, and $\kappa$ is a hyperparameter.
The forward process is formulated as follows:
$p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0})=\mathcal{N}(\bm{x}_{t-1}:\bm{\mu}_{\theta}(\bm{x}_{t},\bm{x}_{0},t),\bm{\Sigma}_{\theta}(\bm{x}_{t},\bm{x}_{0},t))$ (10)
where $\bm{\mu}_{\theta}$ and $\bm{\Sigma}$ are calculated with the trained deep learning model $\epsilon_{\bm{\theta}}$:
$\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},\bm{x}_{0},t)=\frac{\eta_{t-1}}{\eta_{t}}\bm{x}_{t}+\frac{\alpha_{t}}{\eta_{t}}\epsilon_{\bm{\theta}}(\bm{x}_{t},\bm{x}_{0},t),$ (11)
$\bm{\Sigma}_{\bm{\theta}}(\bm{x}_{t},\bm{x}_{0},t)=\kappa^{2}\frac{\eta_{t-1}}{\eta_{t}}\alpha_{t}\textbf{I}$ (12)
Figure 5 demonstrates that the proposed super-resolution method achieves results comparable to diffusion-based adversarial purification [12], which do not require the calculation of gradient.
IV-B3 Adversarial Purification
$y=\{f(\text{SR}(\text{AA}(\bm{x}_{\text{adv}}))))\}$ (13)
IV-C Comparisons to DiffPure
From the perspective of diffusion models, our adversarial super-resolution approach is similar to DiffPure [11], as both utilize backward processes of diffusion models for purification starting from adversarial images as input. However, DiffPure does not employ the entire diffusion process for purification because the diffusion model of DiffPure focuses on the image generation task and begins with Gaussian noise, which can result in the loss of label semantics when using the full diffusion process. Consequently, their defense performance is inferior to methods that use the complete diffusion process, such as MimicDiffusion [12].
Our adversarial super-resolution can leverage the entire diffusion process because it starts with a low-resolution image. This approach overcomes the limitations of DiffPure, which uses only a partial diffusion process. Additionally, super-resolution tasks ensure that the purified image remains consistent with the input image by relying on its low-resolution input, rather than using gradient guidance. This method is more computationally efficient than guidance-based purification, such as MimicDiffusion.
Another reason super-resolution yields better performance is due to the training process. The super-resolution diffusion model in our study is trained using low-resolution images generated by blurring with a filter and adding Gaussian noise. The blurring with a filter helps learn prior knowledge from clean data, similar to our proposed adversarial anti-aliasing. The addition of Gaussian noise can be seen as introducing adversarial perturbations to the clean images. This prior knowledge enables super-resolution to be effectively used for adversarial purification. Furthermore, it suggests that training the diffusion model with adversarial priors can enhance the performance of purification.
We give a trivial theoretical analysis of why adversarial super-resolution tasks achieve better performance than DiffPure’s purification. In DiffPure’s paper, it is proved that for clean distribution $\bm{x}_{t}\sim p(x)$ and adversarial distribution $\bm{\hat{x}}_{t}\sim q(x)$ at diffusion timestep $t$, we have:
$\frac{\partial D_{KL}(\bm{x}_{t}||\bm{\hat{x}}_{t})}{\partial t}\leq 0$ (14)
where the equality happens only when $x_{t}$ and $\hat{x}_{t}$ are sampled from the same distribution, i.e., from the initial timestep $T$. In other words, the difference in the distribution of $p(x)$ and $q(x)$ increases during the reverse diffusion process.
Therefore, it is trivial that utilizing the whole diffusion process from timestep $T$ achieves better performance to recover $\bm{x}_{0}$ from $p(\bm{x})$ than starting from an intermediate timestep $t^{*}$, where:
$D_{KL}(\bm{x}_{T}||\bm{\hat{x}}_{T})<D_{KL}(\bm{x}_{t}^{*}||\bm{\hat{x}}_{t}^{*})$ (15)
Equation 15 demonstrates that our adversarial purification method is superior to DiffPure, which starts from $\hat{x}_{T}$ for purification while DiffPure relies on intermediate timestep $\hat{x}_{t}^{*}$ for purification. However, as ResShift uses only 15 steps for super-resolution, it is important to use adversarial anti-aliasing before the adversarial super-resolution for the reason ResShift achieves fewer diffusion-based purification steps. Compared to guidance-based purification like MimicDiffusion, which also uses the whole diffusion timestep for purification, our method is significantly time-efficient without using a gradient of the diffusion model to maintain the consistency between the purified image and the adversarial image.
V Gradient-Free Adversarial Deblurring Fine-Tuning
In Section IV, we employ the pre-trained ResShift checkpoint for adversarial purification. Intuitively, the performance of ResShift could be further improved by fine-tuning it on the adversarial deblurring task using the target dataset, whereas the original ResShift model is trained solely on the ImageNet dataset.
V-A Deblurring Fine-tuning
Unlike the super-resolution task, which reconstructs a high-resolution image from a sub-sampled low-resolution input, the deblurring task is more suitable for adversarial purification, as it aims to recover the original image from its blurred counterpart while preserving the same resolution. For fine-tuning, we adopt our adversarial anti-aliasing operator as the deblurring kernel. The fine-tuning loss is defined in accordance with ResShift:
$\mathcal{L}(\bm{x}_{t},\bm{x}_{0},t)=\sum\nolimits_{t}\|\hat{\bm{x}}^{t}_{0}-\bm{X}\|_{2}^{2}+\lambda l_{p}(\hat{\bm{x}}^{t}_{0}-\bm{X}),\bm{x}_{0}=\text{AA}(\bm{X})$ (16)
where $\hat{\bm{x}}^{t}_{0}$ is the prediction of $\bm{x}_{0}$ at timestep $t$ by $\epsilon_{\theta}$ and $l_{p}$ is the LPIPS metric.
V-B Contrastive Learning-based Adversarial Deblurring Fine-tuning
To integrate the adversarial purification objective, we incorporate contrastive learning into the deblurring fine-tuning process. Recent studies have shown that contrastive learning achieves superior performance in unsupervised representation learning tasks by pulling representations of positive pairs closer while pushing those of negative pairs farther apart. In this work, we adopt the InfoNCE [41] contrastive objective for fine-tuning. Given a blurred image $\bm{x}_{0}=\text{AA}(\bm{X})$, we construct the positive sample $\bm{x}_{0}^{+}$ as $\bm{x}_{0}^{+}=\text{AA}(\bm{X}+\bm{\varepsilon}),\bm{\varepsilon}\sim\mathcal{N}(0,\mathbf{I})$ where the positive sample is generated by adding Gaussian noise to the high-resolution image $\bm{X}$. For negative pairs $\bm{x}_{0}^{-}$, we employ adversarial examples generated by PGD and AutoAttack with $\ell_{\infty}=4/255$ and $\ell_{2}=4/255$, thereby incorporating adversarial deblurring into the fine-tuning task. The final learning objective is formulated as:
$\mathcal{L}_{\text{CL}}=-\log\frac{\exp(\text{sim}(\bm{x}_{t},\bm{x}_{t}^{+})/\tau)}{\sum_{n\in(\bm{x}_{t}^{+},\bm{x}_{t}^{-})}\exp(\text{sim}(\bm{x}_{t},\bm{x}_{t}^{n})/\tau)}$ (17)
where $\tau$ is the temperature parameter, and sim is the cosine similarity from the $\epsilon_{\theta}$ in feature space:
$\text{sim}(\bm{x},\bm{y})=\frac{\epsilon_{\theta}(\bm{x})\cdot\epsilon_{\theta}(\bm{y})}{\|\epsilon_{\theta}(\bm{x})\|-\|\epsilon_{\theta}(\bm{y})\|}$ (18)
The detailed fine-tuning procedure is summarized in Algorithm 1. To mitigate overfitting, we inject Gaussian noise into the training data.
Algorithm 1 Adversarial Deblurring Fine-tuning
1:Target dataset $\mathcal{D}$, Pre-trained model $\epsilon_{\theta}$, Pre-trained ResNet-50 model $f(\cdot)$
2:repeat
3:  $\bm{X}\sim\mathcal{T}$, $\bm{x}_{0}=\text{AA}(\bm{X}+\bm{\varepsilon}),~\bm{\varepsilon}\sim\mathcal{N}(0,\mathbf{I})$
4:  Construct positive sample
5:  $\bm{x}_{0}^{+}=\text{AA}(\bm{X}+\bm{\varepsilon}),\bm{\varepsilon}\sim\mathcal{N}(0,\mathbf{I})$
6:  Construct negative sample against $f(\cdot)$
7:  $\bm{x}_{0}^{-}=\text{AA}(\text{Adv}(\bm{X})),\text{Adv}\in{\{\text{PGD},\text{AutoAttack}\}},|\text{Adv}|=\{\ell_{\infty}:4/255,\ell_{\text{2}}:4/255\}$
8:  $t\sim\text{Uniform}\left(\{1,\cdots,T\}\right)$
9:  $\bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0},\bm{y}_{0})$
10:  Calculate $\mathcal{L}$ with Equation 16
11:  $\bm{x}_{t}^{+}\sim q(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{0}^{+})$$\bm{x}_{t}^{-}\sim q(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{0}^{-})$
12:  Calculate $\mathcal{L}_{\text{CL}}$ with Equation 17
13:  Fine-tuning $\epsilon_{\theta}$ with $\nabla({\mathcal{L}+}\mathcal{L}_{\text{CL}})$
14:until converged
VI Experiments
VI-A Experimental Setup
@@
TABLE I: The standard and robust accuracy against left: AutoAttack ($\ell_{\infty}=8/255$), right: PGD+EOT ($\ell_{\infty}=8/255$) on the CIFAR-10 dataset. The PGD+EOT is performed in white-box attack scenario with the surrogate process.
Method Target Model Standard Acc(%) AutoAttack Acc(%) PGD+EOT Acc(%)
@@
[…]
TABLE II: The defense performance against AutoAttack ($\ell_{\text{2}}=0.5$) on the CIFAR-10 dataset.
Method Target Model Standard Accuracy(%) Robust Accuracy(%)
Rony et al. [42] WideResNet-28-10 89.05 66.41
Ding et al. [43] WideResNet-28-10 88.02 67.77
Rebuffi et al. [21] WideResNet-28-10 87.50 78.32
Nie et al. [11] WideResNet-28-10 89.23 78.98
Song et al. [12] WideResNet-28-10 92.10 81.52
Ours WideResNet-28-10 92.54 $\pm$ 1.66 84.90 $\pm$ 2.82
[…]
Rebuffi et al. [21] WideResNet-70-16 88.54 80.86
Gowal et al. [6] WideResNet-70-16 88.74 74.03
Nie et al. [11] WideResNet-70-16 91.04 81.17
Song et al. [12] WideResNet-70-16 93.25 83.60
Ours WideResNet-70-16 93.42 $\pm$ 1.51 87.60 $\pm$ 2.35
[…]
TABLE III: The defense performance against C&W ($\ell_{\text{2}}=8/255$, EOT=50) on the CIFAR-10 dataset.
Method Target Model Standard Accuracy(%) Robust Accuracy(%)
Nie et al. [11] WideResNet-28-10 89.23 47.65
Wang et al. [32] WideResNet-28-10 84.85 21.71
Song et al. [12] WideResNet-28-10 92.10 89.91
Ours WideResNet-28-10 92.54 $\pm$ 1.66 90.51 $\pm$ 4.52
[…]
@@
TABLE IV: The standard and robust accuracy against BPDA ($\ell_{\infty}=8/255$) on the CIFAR-10 dataset with WideResNet-28-10 as the target model.
@@
Wang et al. [32] Diffusion 84.45 77.31
Chen et al. [13] Diffusion 91.46 64.73
Song et al. [12] Diffusion 92.10 76.45
Ours Diffusion 92.54 $\pm$ 1.66 81.24 $\pm$ 2.51
[…]
TABLE V: The standard and robust accuracy against AdvDiff on the CIFAR-10 dataset.
@@
[…]
Dataset and target models. We consider CIFAR-10 [44] and ImageNet [45] for major evaluation. The CIFAR-10 dataset consists of 32 × 32 color images across 10 classes, with 50,000 training samples and 10,000 test samples. The ImageNet subset contains 256 × 256 color images from 1000 classes with 1.2 million data in total. For target models, we adopt WideResNet-28-10 and WideResNet-70-16 [46] for CIFAR-10 dataset and ResNet50 [1] for ImageNet dataset. Vision transformers are further included in the experiment on ImageNet dataset. These are commonly adopted backbones for adversarial robustness evaluation.
Attack settings. We evaluate our method with both perturbation-based attacks and diffusion-based unrestricted adversarial attacks. For perturbation-based attacks, we select AutoAttack [48] and PGD [4]. For diffusion-based unrestricted adversarial attacks, we use DiffAttack [49] and AdvDiff [15] for comparisons. DiffAttack is only evaluated on the ImageNet dataset according to the original paper. To ensure a fair comparison with previous diffusion-based adversarial purification, we include the evaluation against the adaptive attack, i.e., reverse pass differentiable approximation (BPDA) [50]. We also evaluate the performance against PGD+EOT that is discussed in [29]. We evaluate adaptive attacks in a white-box setting, where the attacker has access to the gradient of the purification model. However, since our defense employs in-place operations that clip gradients in the adversarial anti-aliasing, we restrict gradient computation to the diffusion-based purification component. On CIFAR-10, the attack settings follow DiffPure [11]. On ImageNet, we randomly sample 5 images from each class and average over 10 runs. The PGD+EOT settings all follow Lee et al. [29] with the surrogate process.
[…]
Evaluation metrics. Following Nie et al. [11], we use standard accuracy and robust accuracy as the evaluation metrics, where standard accuracy measures the clean accuracy against the benign data after the purification and robust accuracy measures the clean accuracy against the adversarial examples after the purification. Both are calculated according to the top-1 classification accuracy. The adversarial examples are generated from the test set.
VI-B Attack Performance
VI-B1 CIFAR-10
Perturbation-based adversarial attack. Table I, II, III presents the defense performance against AutoAttack, PGD, and C&W attacks on the CIFAR-10 dataset. The results demonstrate that our proposed method achieves better standard accuracy and robust accuracy than previous attack methods with our baseline defense. By fine-tuning over the CIFAR-10 dataset, our defense can further improve the defense performance, which validate the effectiveness of the proposed fine-tuning approach. Because images in the CIFAR-10 dataset are only with $32\times 32$ resolution, we set our anti-aliasing filter to a relatively small size. Table IV indicates that the robustness performance of the proposed method is on par with the state-of-the-art method [11]. This finding suggests that our method is more suitable for high-resolution images, as $32\times 32$ may not be large enough to effectively extract the semantic shape for our approach. However, we can further enhance our performance by incorporating adversarial purification techniques from previous work.
Our defense’s performance against PGD+EOT and BDPA on Table I and IV showcases its ability to defend against adaptive attacks. This is because our approach focuses on extracting and recovering the semantic features from adversarial images, rather than inferring and denoising the adversarial perturbations. Moreover, the proposed adversarial anti-aliasing employs in-place operations to block gradient propagation to the attacker, thereby significantly enhancing the robust accuracy of diffusion-based purification under white-box attack scenarios. Consequently, our defense achieves comparable effectiveness against both adaptive and standard attacks. We further suggest that existing diffusion-based purification methods can incorporate our pre-processing approach to better maintain robustness against adaptive attacks.
TABLE VI: The standard and robust accuracy against left: PGD ($\ell_{\infty}=4/255$), right: PGD+EOT ($\ell_{\infty}=4/255$) on the ImageNet dataset. The PGD+EOT is performed in white-box attack scenario with the surrogate process.